### v3

@@ 3 Mapping multinational ownership networks
We construct a global dataset of multinational enterprises (MNEs) and their ownership structures using firm-level data from the Orbis database [^5] The Orbis ownership database, a product of Moody’s, has been widely used in research on MNEs. See, for example, Fonseca et al. (2023), Del Prete and Rungi (2017), and Cravino and Levchenko (2016). Additional country-level covariates complement our analyses as needed, and they are described in Appendix Table C1.
When we refer to ownership structures, we mean the hierarchical control perimeter established by a parent company over its affiliates, which may be organized in ownership chains. The hypothetical chain of ownership in Figure 2 illustrates this concept. A subsidiary is directly controlled if the parent company holds an absolute majority of the voting rights in its shareholders’ meeting (e.g. parent company directly controls subsidiary A in Figure 2). Indirect control occurs when a subsidiary is owned through an intermediate affiliate that is itself controlled by the parent (e.g. Parent indirectly controls Subsidiaries B and C through an intermediary in Figure 2).
@@
Our final dataset includes 226,931 parent firms active in 2019 that collectively control 1,785,493 subsidiaries in 190 countries. Table 1 provides a geographic breakdown of parent companies and subsidiaries by host economy. The European Union (EU) accounts for 42% of parent enterprises and 33% of subsidiaries, reflecting the region’s high degree of economic integration and geographical proximity. Within the EU, the Netherlands, Germany, Italy, France and Spain host the largest number of both parent enterprises and subsidiaries. Asia is the second most important region with 14.7% of parent enterprises and 19.3% of subsidiaries. Among the Asian economies, China and Japan account for the largest shares of MNEs in our data set.
@@
Turning to intermediate affiliates, shown in the rightmost columns of the table 1, we find that they are present in all economies, with no country hosting exclusively directly controlled affiliates. In general, their distribution closely follows that of parent enterprises. Two notable exceptions emerge, however. The United States and Australia host subsidiaries that are typically positioned within longer ownership chains, as indicated by their high ratios of intermediate affiliates to parent firms (1.78 and 1.62, respectively). This suggests that MNEs based in these countries rely more heavily on third countries and intermediate affiliates to structure their global operations, probably due to their geographical remoteness. In contrast, Caribbean countries have shorter ownership chains, with the lowest ratio of intermediate affiliates to parent companies (0.22) in our dataset. This suggests that firms operating there tend to minimize the number of intermediate layers in their corporate networks.
In the following sections, we further investigate the role of tax avoidance strategies observed in the Caribbean. Meanwhile, the reliance on longer chains in the US and Australia supports the hypothesis that geographic distance increases the use of intermediate affiliates, a finding consistent with our later empirical analysis.
3.1 Stylized facts I: heterogeneity and relevance
A key observation is that most subsidiaries of MNEs are controlled indirectly through ownership chains. Table 2 classifies all subsidiaries in our dataset based on the type of control exercised by the parent companies, regardless of geographic location or firm size. We find that 54% of subsidiaries are indirectly controlled, underscoring the prevalence of multi-tiered corporate structures. Referring to the hypothetical ownership chain in Figure 2, directly controlled subsidiaries maintain a direct link to the parent company, while indirectly controlled subsidiaries are linked through other subsidiaries, forming multilevel ownership chains.
Table 2: Direct and indirect control of subsidiaries
Type of control N. subsidiaries %
Direct control 827,516 45.81
Indirect control 978,952 54.19
Total 1,806,468 100.00
Complex multinational enterprises that operate through ownership chains account for a disproportionate share of global economic activity, accounting for 95% of total MNE revenues, despite representing only 27% of the firms in our dataset. In Figure 3, we distinguish between MNEs that exclusively use direct control over subsidiaries (left panel) and MNEs that include at least one ownership chain (right panel). While the latter group is numerically smaller, it dominates global sales, reinforcing the idea that larger corporate networks tend to be structured through multi-tiered ownership chains. [^7] Our findings are consistent with previous studies by Rungi et al. (2017); Altomonte and Rungi (2013), which show that the largest percentile of MNEs by corporate perimeter accounts for approximately 75% of global sales. This pattern emerges because MNEs without ownership chains tend to be smaller in size. On average, they generate 0.107 billion USD in revenue, compared with 1.782 billion USD for complex MNEs with chains of ownership. [^8] Ownership data tend to be more comprehensive than financial accounts because tax authorities require full disclosure of shareholding structures, while financial reporting requirements vary across jurisdictions. As a result, financial data may be incomplete due to exemptions and reporting discrepancies. In our analysis, we assume that missing sales data are not systematically correlated with firm size. Alternative imputation methods do not significantly alter our results, as confirmed in Table 3. Appendix Figure C1 illustrates that sample selection bias is often driven by geographic coverage rather than firm size.
Figure 3: MNEs, ownership chains, and their economic weight
MNEs exhibit considerable heterogeneity in their organizational structure, varying in the number of subsidiaries, ownership chains, host countries, and industries in which they operate. In Figure 4, we classify MNEs according to the number of subsidiaries they control (x-axis) and report their frequency distribution (y-axis). The first bar represents the simplest MNEs, which control only one subsidiary. More complex MNEs, owning at least two subsidiaries, can choose to structure them along an ownership chain. The distribution is highly skewed, with a long right tail, indicating that only 1.2% of MNEs control more than 100 subsidiaries. This heterogeneity goes beyond the number of subsidiaries controlled. MNEs also vary significantly in the number of countries in which they operate and the diversity of industries in which their subsidiaries operate, further highlighting the complexity of global corporate structures.
Figure 4: MNEs’ corporate perimeters
Table 3 provides further evidence of the heterogeneity of MNEs’ corporate structures. Looking first at the number of subsidiaries controlled, we find an average of 7.93 subsidiaries per parent company. But this masks a highly skewed distribution. The median MNE controls only one subsidiary, while the third quartile of MNEs reaches three subsidiaries per parent company. Moving to the right tail, we find that MNEs at the 90th percentile have more than 10 subsidiaries, and only at the 99th percentile do we reach the most complex MNEs, which control more than 120 subsidiaries, with a maximum of 5,075 subsidiaries worldwide. [^9] This is the case of Berkshire Hathaway, the largest multinational conglomerate holding company in the United States, ultimately owned by Warren Buffett.  [^10] In Appendix Figure C2, we estimate the distribution of controlled subsidiaries using a negative binomial regression, which yields a mean of 1.882 and an overdispersion parameter of 0.581. The figure visually represents a log-log plot of the frequency distribution of subsidiaries per parent. A similarly skewed pattern is observed when focusing on intermediate affiliates. The median MNE has zero intermediate affiliate, but the right tail expands sharply, with a jump from 2 to 20 subsidiaries between the 90th and 99th percentiles.
Looking at geographic dispersion, Table 3 shows that the average MNE operates in 2.81 countries. Since our definition of an MNE requires at least one foreign subsidiary, the minimum number of countries is two. However, the 90th percentile includes firms operating in four countries, while at the 99th percentile the number jumps to 19 countries and reaches a maximum of 147.111111This extreme case corresponds to DHL, the globally recognized logistics and courier group. Finally, industry diversification follows a similar pattern. The most complex MNEs, those in the right tail of the distribution, operate in more than 15 industries at the 99th percentile, reaching a maximum of 64 industries. [^12] According to our data, this is the case of HSBC Holdings PLC, originally founded as the Hongkong and Shanghai Banking Corporation and now a British banking group headquartered in London.
Table 3: Heterogeneous distributions of MNEs
Variable Mean st. dev. Min p50 p75 p90 p99 Max
N. controlled subsidiaries 7.93 50.09 1 1 3 10 120 5,075
N. middlemen 1.18 7.72 0 0 1 2 20 688
N. of industries 2.19 3.01 1 1 2 4 15 64
N. of countries 2.81 3.92 2 2 2 4 19 147
The relationship between the number of subsidiaries controlled by an MNE and the maximum length of its ownership chains is illustrated in Figure 5. The plot represents the predicted ranges of a regression model in which the dependent variable is the logarithm of the firm’s scope and the maximum chain length (also in logs) is the key explanatory variable. The relationship appears to be non-linear, with an estimated elasticity of 2.01, suggesting that ownership chains tend to lengthen disproportionately as firms grow. However, large MNEs do not always adopt deep hierarchical structures. Some firms maintain flat ownership networks in which all subsidiaries remain directly controlled by the parent company. [^13] For example, Assyce Fotovoltaica SL, a Spanish provider of solar energy equipment, directly controls approximately 300 subsidiaries without intermediate affiliates.
Figure 5: Number of controlled subsidiaries and maximum ownership chain length
3.2 Stylized facts II: hierarchical positions and geographical distances
So far, we have presented key stylized facts describing MNEs and their ownership chains. In this section, we want to highlight two key patterns linking ownership and geography that underlie the analysis presented in the remainder of this paper. We begin with a measure of the length of ownership chains.
  • On average, an ownership chain consists of three subsidiaries, with 80% of subsidiaries positioned within three hierarchical layers of the parent company. This implies that, in most cases, two intermediate affiliates separate the parent from the final subsidiary.
@@ 3.2 Stylized facts II: hierarchical positions and geographical distances
A defining feature of the organizational complexity of multinational firms is their vertical structure, in which subsidiaries are arranged at multiple levels of the hierarchy, allowing the parent company to coordinate management decisions. This implies that ownership chains can extend over a number of levels and shape the way in which firms manage their global operations.
@@
  • Note: The table reports descriptive evidence on foreign subsidiaries controlled by multinational enterprises. The simplest MNEs with only one direct subsidiary are excluded. Columns 1, 2 and 3 on the left indicate the position in the ownership hierarchy, i.e. how many control links run from the HQs to the single subsidiary. Columns 4, 5, and 6, in the centre, count how many national borders are crossed along an ownership chain to reach a final subsidiary from the headquarters. Columns 7, 8, and 9 on the right report the geographic distance (thousands km) separating final subsidiaries from parent companies. Distance is measured between the most populated cities of hosting countries.
@@
To illustrate these patterns, we present the geography of the ownership chain in Table 4, with domestic affiliates listed separately at the bottom. Columns 1 to 3 report the hierarchical distance of foreign affiliates from their parents, i.e. the number of ownership links required to reach a subsidiary from the headquarters. While a relative majority (39%) of foreign subsidiaries are directly controlled by the parent enterprise (hierarchical level 1, with no intermediate enterprises), the majority are located further down the hierarchy, at level 2 or above. However, as ownership chains lengthen, longer structures become increasingly rare, with only a small fraction of affiliates beyond the tenth hierarchical level. [^14] The longest hierarchical distance observed in our dataset is 21 levels, indicating an ownership chain where 20 intermediate affiliates separate the headquarters from the final subsidiary.
@@
Beyond hierarchical structure, we also examine geographic distance along ownership chains, considering two complementary distance measures. Columns 4 to 6 capture the number of national borders crossed before reaching a subsidiary. Our findings show that most foreign subsidiaries (53%) are located only one country away from their parent. Nevertheless, 37% of the final subsidiaries are integrated into ownership chains that span at least two countries, while 7.6% are located three countries away. Only in exceptional cases do parent firms establish ownership chains that extend up to seven countries (0.02%) from the parent’s home country.
@@
Finally, we take a more classical approach by measuring the physical distance in kilometers between headquarters and subsidiaries, as reported in columns 7 to 9 of Table 4. These data, obtained from Conte et al. (2022) (CEPII database), represent the distance between the main cities in the countries where the headquarters and subsidiaries are located. One could hardly expect that 47.3% of the subsidiaries are located within 5,000 km of the parent company, a somewhat narrower range than expected. Beyond this threshold, the distribution has a long right tail, with subsidiaries located at increasing distances from the parent. [^15] The longest observed distance in our dataset is 19,950 km, between Paraguay and Taiwan.
@@
We illustrate these patterns in Figures 6 and 7, where we analyze the relationship between the affiliate’s position in the ownership chain and its geographic distance from the parent. The relationship is estimated by means of the following equation:
${Y_{p(i)s(j)}=\beta_{0}+\beta^{l}hier_{p(i)s(j)}+\beta_{1}X_{s(j)}++\delta_{p(i)}+\epsilon_{p(i)s(j}}$ (1)
@@
where $Y$ represents either the geographical distance (in thousands of km) or the number of overlapping working hours between the parent in country $i$ and its affiliate in country $j$. The main coefficient of interest, $\beta^{l}$, corresponds to a categorical variable measuring the hierarchical distance ($l$), i.e. the number of ownership links separating the subsidiary from the parent. The model also includes affiliate-level controls such as size and age ($X_{s(j)}$), while parent-level fixed effects ($\delta_{p(i)}$) allow us to capture variation within MNEs.
@@
In Figure 6, we present the estimated coefficients ($\beta^{l}$) with confidence intervals, showing the relationship between hierarchical position and geographic distance. We find a positively increasing and statistically significant distance premium relative to the base category (first-tier subsidiaries). In other words, subsidiaries positioned further down the hierarchy tend to be geographically further away from the parent company. However, the degree of variation is not negligible, as the confidence intervals widen at lower hierarchical levels where there are fewer affiliates.
A complementary pattern emerges in Figure 7, where we analyze the association between hierarchical position and overlapping working hours between the parent and its subsidiaries. Results suggest that as subsidiaries move further down the ownership chain, the number of overlapping hours with the parent systematically decreases. This has critical implications for managerial coordination: subsidiaries at the bottom of longer ownership chains are more likely to operate outside the working hours of parent managers, implying that corporate decisions may need to be delegated to intermediate subsidiaries.
This finding is central to our structural model presented in Section 4. The core intuition is that parent firms structure ownership chains to facilitate efficient managerial coordination. When subsidiaries are geographically distant, direct supervision becomes impractical, necessitating delegation. If this hypothesis holds, then the observed structure of ownership chains reflects the communication needs of managers at different hierarchical levels, providing a rationale for the relationships documented in Figures 6 and 7.
Figure 6: Geographical distance vs hierarchical distance between a parent and its affiliates
@@
  • Note: The figure presents the estimated coefficients from an econometric analysis regressing the geographical distance between parent companies and their affiliates on a categorical variable representing hierarchical distance. A parent-company fixed effect is included to capture within-MNE variation, along with controls for the size and age of intermediate affiliates. The base category corresponds to hierarchical distance 1, where the parent directly controls the subsidiary. Consequently, all coefficients represent the variation in geographical distance relative to this base category. Hierarchical distances of 9 or greater are grouped into a single category. Standard errors, clustered at the parent-company level, are reported in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
@@
Figure 7: Overlapping working hours and hierarchical distance of a parent and its affiliates
@@
  • Note: The figure presents the estimated coefficients from an econometric analysis regressing overlapping working hours—used as a proxy for communication frictions—on a categorical variable representing hierarchical distance between parent companies and their subsidiaries. A parent-company fixed effect is included to capture within-MNE variation, along with controls for subsidiary size and age. The base category corresponds to hierarchical distance 1, where the parent directly controls the subsidiary. Consequently, all coefficients represent the variation in overlapping working hours relative to this base category. Hierarchical distances of 9 or greater are grouped into a single category. Standard errors, clustered at the parent-company level, are reported in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
Finally, in Appendix B, we provide additional evidence using naive unstructured models to analyze the relationship between MNEs’ investment decisions and geography. This exercise also highlights the empirical challenges that researchers and policymakers may face when network information about ownership chains is not taken into account.
4 A structural model for ownership chains

### v4

@@ 3 Mapping multinational ownership networks
We construct a global dataset of multinational enterprises (MNEs) and their ownership structures using firm-level data from the Orbis database [^4] The Orbis ownership database, a product of Moody’s, has been widely used in research on MNEs. See, for example, Fonseca et al. (2023), Del Prete and Rungi (2017), and Cravino and Levchenko (2016).. Additional country-level covariates complement our analyses as needed, and they are described in Appendix Table B1.
When we refer to ownership structures, we mean the hierarchical control perimeter established by a parent company over its affiliates, which may be organized in ownership chains. The hypothetical chain of ownership in Figure 2 illustrates this concept. A subsidiary is directly controlled if the parent company holds an absolute majority of the voting rights in its shareholders’ meeting (e.g. parent company directly controls subsidiary A in Figure 2). Indirect control occurs when a subsidiary is owned through an intermediate affiliate that is itself controlled by the parent (e.g., Parent indirectly controls Subsidiaries B and C through an intermediary in Figure 2).
@@
Our final dataset includes 226,931 parent firms active in 2019 that collectively control 1,785,493 subsidiaries in 190 countries. Indirect control is pervasive among multinational enterprises: 54% of foreign subsidiaries are controlled through at least one intermediate affiliate. At the same time, multinational groups characterized by complex ownership structures account for approximately 95% of global sales by multinationals, underscoring their economic relevance [^6] Appendix Appendix C: Additional stylized facts on multinational enterprises and their corporate structures provides additional details on this evidence, as well as further descriptive statistics on the control networks of multinational enterprises in our sample..
@@
Table 1 provides a geographic breakdown of parent companies and subsidiaries by hosting economy. The European Union (EU) accounts for 42% of parent enterprises and 33% of subsidiaries, reflecting the region’s high degree of economic integration and geographical proximity. Within the EU, the Netherlands, Germany, Italy, France, and Spain host the largest number of both parent enterprises and subsidiaries. Asia is the second most important region with 14.7% of parent enterprises and 19.3% of subsidiaries. Among the Asian economies, China and Japan account for the largest shares of MNEs in our data set.
@@
Turning to intermediate affiliates, shown in the rightmost columns of the table 1, we find that they are present in all economies, with no country hosting exclusively directly controlled affiliates. Generally, their distribution closely follows that of their parent enterprises. Two notable exceptions emerge, however. The United States and Australia host subsidiaries that are typically positioned within longer ownership chains, as indicated by their high ratios of intermediate affiliates to parent firms (2.10 and 2.17, respectively). This suggests that MNEs based in these countries rely more heavily on third countries and intermediate affiliates to structure their global operations, probably due to their geographical remoteness.
The reliance on longer chains in the United States and Australia is consistent with the finding that geographic distance is associated with the use of intermediate affiliates, a pattern also observed in the subsequent empirical analysis. The following sections present stylized facts along these lines, preceded by brief remarks on tax optimization.
3.1 Whither tax optimization?
In this work, we investigate why one or more middlemen exist between a parent company and its subsidiaries. Such ownership configurations are commonly associated with tax optimization strategies and, thus, profit shifting. Specifically, refer to the work by Francois and Vicard (2025) using the same data. They find that the number of layers of ownership between subsidiaries and the parent company correlates with a higher propensity to report zero profits. Moreover, MNEs with sufficiently complex ownership structures are more likely to shift profits away from their high-tax affiliates.
In the following paragraphs, we have a different perspective. We do not challenge the claim that profit shifting is more likely if ownership structures are complex. Yet, we believe that tax avoidance is not the main reason for the origin of such complex ownership structures. Indeed, if the locations of intermediary companies along an ownership chain were primarily driven by tax minimization, one would expect them to differ systematically from those of final subsidiaries, for instance, by being located in low-tax jurisdictions and exhibiting few signs of real economic activity. We estimate the likelihood that a middleman, compared to other subsidiaries, exhibits features typically associated with potential tax evasion practices. The traits we test include the strategic location in a tax haven and hallmarks of a Special Purpose Entity (SPE), such as the absence of employees and non-financial assets, established to obtain jurisdiction-specific advantages.
Column (1) of Table 2 reports predictions at the margin after a logistic regression in which the dependent variable is a dummy for tax-haven location, based on the classification of Hines Jr (2010). The main explanatory variable is a dummy indicating middleman versus final subsidiaries. Firm age and size are included as controls, along with sector fixed effects.
Table 2: Middlemen, tax havens, and SPEs
Predicted probability of being:
(1) (2)
located in a tax haven a suspected SPE
Final subsidiary 0.1445*** 0.0042***
(0.0016) (0.0001)
Middleman subsidiary 0.1628*** 0.0054***
(0.0017) (0.0002)
Difference between predicted probabilities 0.0183*** 0.0012***
(0.0015) (0.0002)
  • Note: The table reports predicted probabilities at the margin after two logistic regressions. In Column (1), the dependent variable is a dummy equal to one if the subsidiary is located in a tax haven; in Column (2), it is a dummy equal to one if the firm exhibits characteristics of a suspected special purpose entity. Predicted probabilities are obtain after estimates of coefficients for a dummy variable indicating whether the subsidiary is a middleman or a final subsidiary. All regressions control for firm size and age, and include sector fixed effects. Standard errors are clustered at the parent level (*** p<<0.01, ** p<<0.05, * p<<0.1).
Column (2) presents predictions at the margin from the same specification, but with a dummy indicating whether the subsidiary is suspected to be an SPE as the outcome. We follow the European Statistical System (ESS) definition of SPE [^7] According to ESS, SPEs are legally established companies with up to five employees, little or no physical presence, and minimal production in the host economy. They are directly or indirectly controlled by foreign entities and transact almost exclusively with nonresidents., also referenced in the IMF guidelines (Harutyunyan et al., 2022) and accordingly define potential SPEs as foreign subsidiaries with up to five employees, no non-financial assets, and no operating revenues.
We find that the predicted probability for a middleman to be located in a tax haven is slightly higher than for a final subsidiary, 16% vs 14%. The difference between the two is statistically significant, albeit minimal [^8] Please note that the location in a country that is classified as a tax haven is not per se the indication that a firm engages in tax avoidance. . Moreover, the likelihood of exhibiting features defining SPEs is very low for both types of subsidiaries. Altogether, results suggest that while tax optimization may partially explain the existence of multiple ownership chains, it accounts for only a small fraction of a broader, more complex reality. In other words, these patterns are not markedly more prevalent among middlemen than other subsidiaries, indicating that the use of middlemen for tax purposes is neither dominant nor exclusive.
We investigate this further by proposing a stylized fact that examines whether there is a correlation between the location along the ownership chain and lower corporate taxes in the hosting country. We propose a basic reduced form as follows:
$Y_{p(i)s(j)}=\beta_{0}+\beta^{l}hier_{p(i)s(j)}+\beta_{1}X_{s(j)}+\epsilon_{p(i)s(j}$} (1)
where $Y$ is the tax burden on a typical subsidiary firm $s$ in the hosting country $j$, which is controlled by a parent $p$ located in a country $i$. Country-level taxation is captured by the WorldBank (2019) total tax and contribution rate (TTCR). The main coefficient of interest, $\beta^{l}$, corresponds to a categorical variable measuring the hierarchical distance ($l$), i.e., the number of ownership links separating the subsidiary from the parent. Please note that we aggregate subsidiaries when they are located in layers above eight. The model also includes affiliate-level controls such as size and age in the matrix $X_{s(j)}$. Correlations are visualized in Figure 3.
Evidently, we do not find any statistically significant pattern along the ownership chain. There is no apparent difference in tax burdens if a subsidiary is directly or indirectly controlled by the parent company. A typical intermediary subsidiary along the ownership chain would pay a similar amount of corporate taxes and mandatory contributions of any other subsidiary. Please note that we cannot confirm the absence of profit shifting or any other tax optimization strategy. What we can say is that the sequence along an ownership chain seems to have a different explanation. Yet, subsidiaries controlled by the same parent company might pursue tax optimization strategies regardless of their position along the ownership chain.
Figure 3: Corporate taxes along ownership chains
  • Note: The figure presents the estimated coefficients of eq. 1, where the level of corporate income taxes in the country hosting subsidiaries is regressed on a categorical variable representing hierarchical distance from the parent company to the affiliate. Controls for the size and age of the subsidiaries are included, together with 2-digit industry fixed effects and parent-level fixed effects. The base category corresponds to hierarchical distance 1, where the parent directly controls the subsidiary. Consequently, all coefficients represent the variation in tax burden relative to this base category. Hierarchical distances above eight are grouped into a single category.
3.2 Ownership chains, distance, and overlapping working hours
In this section, we highlight three key patterns that link ownership and geography. We begin by measuring the lengths of ownership chains.
  • Ownership chains can be long, and they can cross multiple country borders. Approximately 61% of subsidiaries are positioned at least two ownership links’ distance from the parent company, 47% cross at least two country borders, and 53% extend approximately more than 10,000 kilometers.
@@ 3.2 Ownership chains, distance, and overlapping working hours
A defining feature of the organizational complexity of multinational firms is their vertical structure, in which subsidiaries are arranged at multiple levels of the hierarchy, allowing the parent company to coordinate management decisions.
@@
  • Note: The table reports descriptive evidence on foreign subsidiaries controlled by multinational enterprises. The simplest MNEs, which have only one direct subsidiary, are excluded. Columns 1, 2 and 3 on the left indicate the position in the ownership hierarchy, i.e. how many control links run from the HQs to the single subsidiary. Columns 4, 5, and 6, in the centre, count how many national borders are crossed along an ownership chain to reach a final subsidiary from the headquarters. Columns 7, 8, and 9 on the right report the geographic distance (thousands km) separating final subsidiaries from parent companies. Distance is measured between the most populated cities of hosting countries.
@@
To illustrate these patterns, we present the geography of the ownership chain in Table 3, with the number of domestic affiliates listed separately at the bottom. Columns 1 to 3 report the hierarchical distance of foreign affiliates from their parents, i.e., the number of ownership links required to reach a subsidiary from the headquarters. While a relative majority (39%) of foreign subsidiaries are directly controlled by the parent enterprise (hierarchical level 1, with no intermediate enterprises), the majority are located further down the hierarchy, at level 2 or above. However, as ownership chains lengthen, longer structures become increasingly rare, with only a small fraction of affiliates beyond the tenth hierarchical level. [^9] The longest hierarchical distance observed in our dataset is 21 levels, indicating an ownership chain where 20 intermediate affiliates separate the headquarters from the final subsidiary.
@@
Beyond hierarchical structure, we also examine geographic distance along ownership chains, considering two complementary measures. Columns 4 to 6 capture the number of national borders crossed before reaching a subsidiary. Our findings show that most foreign subsidiaries (53%) are located only one country away from their parent. Nevertheless, 37% of the final subsidiaries are integrated into ownership chains that span at least two countries, while 7.6% are located three countries away. Only in exceptional cases do parent firms establish ownership chains that extend up to seven countries (0.02%) from the parent’s home country.
@@
Finally, we take a more classical approach by measuring the physical distance in kilometers between headquarters and subsidiaries, as reported in columns 7 to 9 of Table 3. These data, obtained from Conte et al. (2022) (CEPII database), represent the distance between the main cities in the countries where the headquarters and subsidiaries are located  [^10] Appendix Table B4 also shows the correlation between hierarchical distance of a subsidiary from the parent and the number of country borders crossed..
One could hardly expect that 47.3% of the subsidiaries are located within 5,000 km of the parent company, a somewhat narrower range than expected. Beyond this threshold, the distribution has a long right tail, with subsidiaries located at increasing distances from the parent. [^11] The longest observed distance in our dataset is 19,950 km, between Paraguay and Taiwan.
@@
We illustrate these patterns in Figures 4 and 5, where we analyze the relationship between the affiliate’s position in the ownership chain and its geographic distance from the parent. The relationship is estimated by means of the equation 1 used in the previous paragraphs, where $Y$ represents either the geographical distance (in thousands of km) or the number of overlapping working hours between the parent in country $i$ and its affiliate in country $j$. The latter is defined as the total number of hours within a standard business day during which the working schedules of two firms’ locations overlap, given their respective local time zones and assuming a ten-hour working day. The main coefficient of interest, $\beta^{l}$, corresponds to a categorical variable measuring the hierarchical distance ($l$), i.e., the number of ownership links separating the subsidiary from the parent. The model also includes affiliate-level controls such as size and age ($X_{s(j)}$), while parent-level fixed effects ($\delta_{p(i)}$) allow us to capture variation within MNEs.
@@
In Figure 4, we present the estimated coefficients ($\beta^{l}$) along with their confidence intervals, illustrating the relationship between hierarchical position and geographic distance. We find a positively increasing and statistically significant distance premium relative to the base category (first-tier subsidiaries). In other words, subsidiaries positioned further down the hierarchy tend to be geographically further away from the parent company. However, the degree of variation is not negligible, as the confidence intervals widen at lower hierarchical levels where there are fewer and fewer affiliates.
  • Subsidiaries situated at more distant hierarchical levels in ownership chains have, on average, less overlapping working hours with the ultimate parent company.
Figure 4: Geographical distance vs hierarchical distance of a parent and its affiliates
@@
  • Note: The figure presents the estimated coefficients after the reduced-form eq. 1, where the geographical distance between parent companies and their affiliates is regressed on a categorical variable representing hierarchical distance. Parent-company fixed effect is included, along with controls for the size and age of intermediary affiliates. The base category corresponds to hierarchical distance 1, where the parent directly controls the subsidiary. Consequently, all coefficients represent the variation in geographical distance relative to this base category. Hierarchical distances above eight are grouped into a single category. Standard errors, clustered at the parent-company level, are reported in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
@@
Figure 5: Overlapping working hours vs hierarchical distance of a parent and its affiliates
@@
  • Note: The figure presents the estimated coefficients after the reduced-form eq. 1, where overlapping working hours—used as a proxy for communication frictions—are regressed on a categorical variable representing hierarchical distance between parent companies and their subsidiaries. A parent-company fixed effect is included, along with controls for subsidiary size and age. The base category corresponds to hierarchical distance 1, where the parent directly controls the subsidiary. Consequently, all coefficients represent the variation in overlapping working hours relative to this base category. Hierarchical distances above eight are grouped into a single category. Standard errors, clustered at the parent-company level, are reported in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
Figure 5 reveals a clear hierarchical gradient in time overlap. As subsidiaries sit lower in the ownership chain, their overlapping working hours with the ultimate parent decline monotonically. This pattern carries sharp implications for coordination: subsidiaries at the tail of long chains are more likely to operate outside the parent’s business hours, making direct managerial oversight costly, and increasing the need for delegation to intermediate entities.
We provide further evidence consistent with this intuition by estimating a naive control gravity equation that isolates the role of overlapping working hours across firm locations. We aggregate ownership links at the country-pair level, distinguishing parent–subsidiary from middleman–subsidiary relationships. As illustrated in Figure 2, links from the parent to both intermediaries A and B and final subsidiary C define parent control, while links from intermediaries A and B to the final subsidiary C capture middleman control. Table 4 reports estimates from the following reduced-form equation:
$N_{ij}=exp(\beta^{wh}wh_{ij}+\bm{\beta^{\prime}}\mathbf{x}_{ij}+\gamma_{i}+\gamma_{j})\epsilon_{ij}$} (2)
where $i$ and $j$ stand for origin and destination country and $N_{ij}$ counts the number of companies in country $j$ controlled by companies located in country $i$. $wh_{ij}$ measures overlapping working hours between $i$ and $j$, while $\mathbf{x}_{ij}$ is a vector of gravity controls we source from the CEPII Gravity dataset (namely, the logarithm of distance measured in km, and indicator variables for the presence of, respectively, contiguous borders, common language, common legal origins, colonial ties and regional trade agreements between countries). $\gamma_{i}$ and $\gamma_{j}$ capture origin and destination country fixed effects.
We find that overlapping working hours are significant only in the control relationships between middlemen and final subsidiaries. If real-time communication is central to organization and coordination, this constraint does not bind the location of subsidiaries relative to the parent. Instead, coordination costs appear to be shifted downstream, with middlemen absorbing the burden of synchronizing operations on behalf of the parent.
Table 4: Two naive control gravity equations for MNEs
Dep. Var. $N_{ij}$= # of companies country $i$ controls in country $j$
Sample
Parent-subsidiary
control links
Middleman-final
control links
N. of overlapping 0.006 0.068**
working hours (0.023) (0.027)
RTA 0.136 0.291*
(0.195) (0.168)
Log distance (km) -0.336*** -0.227***
(0.072) (0.088)
Home 2.516*** 3.472***
(0.315) (0.250)
Language 0.747*** 0.641***
(0.089) (0.108)
Colony dependence 0.312** 0.385***
(0.129) (0.124)
Legal origins 0.238*** 0.268***
(0.087) (0.101)
Observations 32,952 35,012
Fixed Effects i,j i,j
  • Notes: $N_{ij}$ counts the number of companies in country $j$ controlled by companies in country $i$. Results in Column (1) refer to the sample of parent-subsidiary control relationships, whereas Column (2) refers to middleman-final control relationships. Standard errors clustered by origin country in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
These findings are central to our structural model presented in Section 4. The core intuition is that parent firms structure ownership chains to facilitate efficient managerial coordination. When subsidiaries are geographically distant, direct supervision becomes impractical, necessitating delegation. If this hypothesis holds, then the observed structure of ownership chains reflects the communication needs of managers at different hierarchical levels, providing a rationale for the relationships documented in Figures 4 and 5.
4 An analytical framework for ownership chains